OpenStudy (anonymous):

this equation 3x^4 - 4x^3 - 2x^2= 0 has three real solutions. can someone please tell me what they are?

OpenStudy (anonymous):

x = 1/3 (2+sqrt(10)) x = 1/3 (2-sqrt(10)) x = 0

OpenStudy (anonymous):

\[x \approx 0, 1.72, -0.39\]

OpenStudy (anonymous):

so: x=0 x=1.72 x=-0.39

OpenStudy (anonymous):

these figures dont work for me

OpenStudy (anonymous):

my software wont accept them

OpenStudy (anonymous):

umm what figures?

OpenStudy (anonymous):

the 0, 1.72, -0.38

OpenStudy (anonymous):

you might have to retype them, copying from this site isnt really the best on other softwares

OpenStudy (anonymous):

And its not right to just copy and paste, its against the code of conduct.

OpenStudy (anonymous):

i actually got the same answer when i did it myself and typed it into my software and it rejected so i came here to see if someone had a different answer

OpenStudy (mathstudent55):

First, factor x^2 from every term.

OpenStudy (anonymous):

hmm, sorry then, stupid response on my part. but if it helps, im getting the same thing, try more digits maybe...?

OpenStudy (mathstudent55):